OpenStudy (anonymous):

if a(4,-15) and b(7,-9), find ab

OpenStudy (anonymous):

What's the distance?

OpenStudy (anonymous):

Distance formula: =√((7-4)^2+(〖(-9)-(-15)〗^2 ) ) =√((7-4)^2+(-9+15)^2 ) √(〖(3)〗^2+〖(6)〗^2 ) =√(9+36) =√45 =∛5 or 6.708203933

OpenStudy (anonymous):